# Double Successive Over-Relaxation Q-Learning with an Extension to Deep Reinforcement Learning

### v1

@@ Abstract
Q-learning is a widely used algorithm in reinforcement learning (RL), but its convergence can be slow, especially when the discount factor is close to one. Successive Over-Relaxation (SOR) Q-learning, which introduces a relaxation factor to speed up convergence, addresses this issue but has two major limitations: In the tabular setting, the relaxation parameter depends on transition probability, making it not entirely model-free, and it suffers from overestimation bias. To overcome these limitations, we propose a sample-based, model-free double SOR Q-learning algorithm. Theoretically and empirically, this algorithm is shown to be less biased than SOR Q-learning. Further, in the tabular setting, the convergence analysis under boundedness assumptions on iterates is discussed. The proposed algorithm is extended to large-scale problems using deep RL. Finally, the tabular version of the proposed algorithm is compared using roulette and grid world environments, while the deep RL version is tested on a maximization bias example and OpenAI Gym environments.
@@ I Introduction
Reinforcement learning (RL) is a commonly employed approach for decision-making in a variety of domains, wherein agents learn from their interactions with the environment. The primary goal of RL is to find an optimal policy, which involves solving Bellman’s optimality equation. Depending on the amount of information available, various algorithms are proposed in both dynamic programming and RL to solve Bellman’s optimality equation. When the complete information of the Markov Decision Process (MDP) is available, techniques from dynamic programming, such as value iteration and policy iteration, are utilized to find the optimal value function, thereby, an optimal policy [1]. In cases where the MDP is not fully known (model-free), RL techniques are utilized [2]. Q-learning, introduced by Watkins in 1992, is a fundamental algorithm in RL used to find an optimal policy in model-free environments [3]. Despite its wide application in several areas (see, [4, 5, 6, 7]), Q-learning suffers from slow convergence and over-estimation bias. The drawbacks of Q-learning are well-documented in the literature, and considerable efforts are being made to overcome them (for instance, see [8, 9, 10, 11, 12]).
@@
In 1973, a successive over-relaxation (SOR) technique was proposed by D. Reetz to fasten the convergence of value iteration [13]. This technique involves constructing a modified Bellman equation with a successive over-relaxation parameter. Recently, this technique has been incorporated into Q-learning, resulting in a new algorithm called SOR Q-learning (SORQL)[10]. The underlying deterministic mapping of the SORQL algorithm has been shown to have a contraction factor less than that of the standard Q-Bellman operator [10]. This was achieved by selecting a suitable relaxation factor that depends on the transition probability of a self-loop of an MDP, making the algorithm partially model-free. Furthermore, the theory and our empirical results on benchmark examples have revealed that SOR Q-learning, similar to Q-learning, suffers from overestimation bias. To address these issues, we propose double SOR Q-learning (DSORQL), which combines the idea of double Q-learning [8] with SOR. The advantages of the proposed algorithm are that it is entirely model-free and has been shown, both theoretically and empirically, to have less bias than SORQL. In addition, this work provides the convergence analysis of the proposed algorithm using stochastic approximation (SA) techniques. Recently, the concept of SOR is applied to the well-known speedy Q-learning algorithm [11], leading to the development of a new algorithm called generalized speedy Q-learning, which is discussed in [14]. Further, in [15], the SORQL algorithm has been suitably modified to handle two-player zero-sum games, and consequently, a generalized minimax Q-learning algorithm was proposed. It is also worth noting that a deep RL version of the SOR Q-learning algorithm, similar to the deep Q-network (DQN) [16], is proposed in [17]. Additionally, the application of this deep RL variant of the SOR Q-learning algorithm is discussed for auto-scaling cloud resources [17].
@@
The convergence of several RL algorithms depends on the theory developed in SA [18, 19, 20, 21]. Many of the iterative schemes in the literature assume the boundedness and proceed to show the convergence. For instance, in [15], the convergence of the generalized minimax Q-learning algorithm was shown to converge under the boundedness assumption. In this manuscript, we take the same approach to show the convergence of the proposed algorithms in the tabular setting. Motivated from the work in [16], and [17], wherein the double Q-learning and SOR Q-learning in the tabular version extended to large-scale problems using the function approximation. In this manuscript, we extend the tabular version of the proposed algorithm and present a double successive over-relaxation deep Q-network (DSORDQN). The proposed algorithm is tested on OpenAI gym’s CartPole [22], LunarLander [22] environments, and a maximization example where the state space is in the order of $10^{9}$ [23]. The experiments corroborate the theoretical findings.
The remainder of the paper is organized as follows: In Section II, we discuss the results required for proving the convergence of the proposed algorithms, along with basic notations and terminologies. Section III presents the proposed algorithms, while Section IV addresses the convergence analysis of the proposed algorithms and provides insights into the theoretical understanding of their bias. Numerical experiments are conducted in Section V to demonstrate the effectiveness of the proposed algorithms. Finally, we conclude the manuscript with Section VI.
@@ II Preliminaries
$r(i,a)=\sum_{j=1}^{|S|}p^{j}_{ia}r^{j}_{ia},\quad\text{and}\quad\mathcal{M}Q(j)=\max_{b\in A}Q(j,b).$
To simplify the equations, we defined the max-operator $\mathcal{M}Q(i)$ as $\max_{a\in A}Q(i,a)$ for $i\in S$. Also, $R_{\max}=\max_{(s,a,s^{\prime})}|r^{s^{\prime}}_{sa}|,\forall(s,a,s^{\prime})\in S\times A\times S$. These notations will be used throughout the manuscript as necessary. Note that $U$ is a contraction operator with $\gamma$ as the contraction factor, and $Q^{*}$ serves as the optimal action-value function of the operator $U$. Moreover, one can obtain an optimal policy $\pi^{*}$ from the $Q^{*}$, using the relation $\pi^{*}(i)\in\text{arg}\,\mathcal{M}Q^{*}(i),\;\forall i\in S.$ Q-learning algorithm is a SA version of Bellman’s operator $U$. Given a sample $\{i,a,j,r^{j}_{ia}\}$, current estimate $Q_{n}(i,a)$, and a suitable step-size $\beta_{n}(i,a)$, the update rule of Q-learning is as follows
$Q_{n+1}(i,a)=Q_{n}(i,a)+\beta_{n}(i,a)(r^{j}_{ia}+\gamma\,\mathcal{M}Q_{n}(j)-Q_{n}(i,a)).$
@@
$=w\bigg{(}r(i,a)+\gamma\sum_{j=1}^{|S|}p^{j}_{ia}\mathcal{M}Q(j)\bigg{)}+(1-w)\mathcal{M}Q(i),$
where $0<w\leq w^{*}$, and $w^{*}=\min\limits_{i,a}\left\{\frac{1}{1-\gamma p^{i}_{ia}}\right\}.$
In [10], it was shown that $U_{w}$ is $1-w+w\gamma$-contractive. Moreover, from Lemma 4, in [10] for $1\leq w\leq w^{*}$ the contraction factor of the operator $U_{w}$ satisfy $1-w+\gamma w\leq\gamma$. Further, it was shown that
$\mathcal{M}Q^{*}_{w}(i)=\mathcal{M}Q^{*}(i),\,\forall i\in S,$ (2)
@@
$\hskip 56.9055pt+(1-w)\mathcal{M}Q_{n}(i)\Big{]},$ (3)
where $0<w\leq w^{*}$, and $0\leq\beta_{n}(i,a)\leq 1$ converges to the fixed point of $U_{w}$ w.p.1 [10]. Note that the successive relaxation factor in the above equation depends on the transition probability, which makes it not entirely model-free. Owing to this and the natural concerns regarding the over-estimation of SORQL. This manuscript proposes a model-free variant of the SORQL algorithm and a model-free double SOR Q-learning algorithm. To prove the convergence of the proposed algorithms in the tabular setting, the following result from [21] will be useful, and we conclude this section by stating it here as lemma.
@@
(Lemma 1, in [21]) Let a random process $\left(\Phi_{n},J_{n},\beta_{n}\right),$ $n\geq 0$, where $\Phi_{n},J_{n},\beta_{n}:X\rightarrow\mathbb{R}$ satisfy the following relation:
@@ III-A Tabular version:
$\widetilde{P}^{j}_{ia}=\dfrac{Y_{n}[i][j][a]}{n},\quad\forall n\geq 1.$
@@
$w_{n+1}=w_{n}+\alpha(n)(f(w_{n})+\epsilon_{n}),\quad n\geq 1,$ (5)
@@
$f(w_{n})=\dfrac{1}{1-\gamma\min\limits_{i,a}p^{j}_{ia}}-w_{n},\;\epsilon_{n}=\dfrac{1}{1-\gamma\min\limits_{i,a}\widetilde{P}^{j}_{ia}}-w^{*}.$
@@
As a result of the above observations and [20], one can obtain the following theorem
@@
7:         $b^{*}=\arg\max_{b}Q^{A}_{n}(i^{\prime}_{n},b)$
8:         $c^{*}=\arg\max_{c}Q^{A}_{n}(i_{n},c)$
9:         $Q^{A}_{n+1}(i_{n},a_{n})=(1-\beta_{n}(i_{n},a_{n}))Q^{A}_{n}(i_{n},a_{n})+\beta_{n}(i_{n},a_{n})\big{(}w_{n}(r^{\prime}_{n}+\gamma Q^{B}_{n}(i^{\prime}_{n},b^{*}))+(1-w_{n})Q^{B}_{n}(i_{n},c^{*})\big{)}$
10:     else
11:         $d^{*}=\arg\max_{d}Q^{A}_{n}(i^{\prime}_{n},d)$
12:         $e^{*}=\arg\max_{e}Q^{A}_{n}(i_{n},e)$
13:         $Q^{B}_{n+1}(i_{n},a_{n})=(1-\beta_{n}(i_{n},a_{n}))Q^{B}_{n}(i_{n},a_{n})+\beta_{n}(i_{n},a_{n})\big{(}w_{n}(r^{\prime}_{n}+\gamma Q^{A}_{n}(i^{\prime}_{n},d^{*}))+(1-w_{n})Q^{A}_{n}(i_{n},e^{*})\big{)}$
14:     end if
15:     $i_{n+1}=i^{\prime}_{n}$
16:end for
17:return $Q_{N-1}$
@@
If the sequence $\{w_{n}\}_{n\geq 0}$ in Algorithm 1 is a constant sequence, with $w$ as the constant and $0<w\leq w^{*}$, we refer to this algorithm as double SOR Q-learning (DSORQL). Therefore, given the samples $\{i,a,j,r^{j}_{ia}\}$ the update rule for the DSORQL is as follows:
With probability $0.5$ update $\widetilde{Q}^{A}$:
$\widetilde{Q}^{A}_{n+1}(i,a)$ $=(1-\beta_{n})\widetilde{Q}^{A}_{n}(i,a)$
@@
else update $\widetilde{Q}^{B}$:
$\widetilde{Q}^{B}_{n+1}(i,a)$ $=(1-\beta_{n})\widetilde{Q}^{B}_{n}(i,a)$
$\;+\beta_{n}\Big{[}w\big{(}r^{j}_{ia}+\gamma\widetilde{Q}^{A}_{n}(j,d^{*})\big{)}+(1-w)\big{(}\widetilde{Q}^{A}_{n}(i,e^{*})\big{)}\Big{]},$
@@ III-B Deep RL version:
The proposed double SOR deep Q-network (DSORDQN) is a deep RL version of the tabular DSORQL. DSORDQN utilizes the technique of SOR method with DDQN. Specifically, similar to DDQN we use online Q-network $Q(S,a;\theta)$ for selecting the actions, and the target Q-network $Q(S,a;\theta^{\prime})$, to evaluate the actions. The gradient descent will be performed on the following loss function: $L(\theta_{i})=E\big{[}(y_{j}-Q(s,a;\theta_{i}))^{2}\big{]},$ where $y_{j}=w(r(i_{j},a_{j})+\gamma Q(i_{j+1},\operatorname*{arg\,max}_{b}Q(i_{j+1},b;\theta_{i});\theta^{\prime}_{i})+(1-w)Q(i_{j},\operatorname*{arg\,max}_{e}Q(i_{j},e;\theta_{i});\theta^{\prime}_{i})$ for sample $(i_{j},a_{j},r(i_{j},a_{j}),i_{j+1})$ from the replay buffer with $w>1$. Interestingly, the advantages of DDQN and SORDQN transfer to the proposed DSORDQN algorithm, resulting in more stable and reliable learning at scale.

### v2

@@ Abstract
Q-learning is a widely used algorithm in reinforcement learning (RL), but its convergence can be slow, especially when the discount factor is close to one. Successive Over-Relaxation (SOR) Q-learning, which introduces a relaxation factor to speed up convergence, addresses this issue but has two major limitations: In the tabular setting, the relaxation parameter depends on transition probability, making it not entirely model-free, and it suffers from overestimation bias. To overcome these limitations, we propose a sample-based, model-free double SOR Q-learning algorithm. Theoretically and empirically, this algorithm is shown to be less biased than SOR Q-learning. Further, in the tabular setting, the convergence analysis under boundedness assumptions on iterates is discussed. The proposed algorithm is extended to large-scale problems using deep RL. Finally, both the tabular version of the proposed algorithm and its deep RL extension is tested on benchmark examples.
@@ I Introduction
Reinforcement learning (RL) is a commonly employed approach for decision-making in a variety of domains, wherein agents learn from their interactions with the environment. The primary goal of RL is to find an optimal policy, which involves solving Bellman’s optimality equation. Depending on the amount of information available, various algorithms are proposed in both dynamic programming and RL to solve Bellman’s optimality equation. When the complete information of the Markov Decision Process (MDP) is available (model-based), techniques from dynamic programming, such as value iteration and policy iteration, are utilized to find the optimal value function, thereby, an optimal policy [1]. In cases where the MDP is not fully known (model-free), RL techniques are utilized [2]. Q-learning (QL), introduced by Watkins in 1992, is a fundamental algorithm in RL used to find an optimal policy in model-free environments [3]. Despite its wide application in several areas [4, 5, 6, 7], Q-learning suffers from slow convergence [8, 9] and over-estimation bias [10, 11, 12, 13].
In 1973, D. Reetz proposed a successive over-relaxation (SOR) technique to accelerate the convergence of value iteration in dynamic programming [14]. This technique involves constructing a modified Bellman optimality equation using an SOR parameter. The SOR parameter in [14] depends on the transition probability of the self-loop of an MDP. Furthermore, it was demonstrated in [14] that, with a proper choice of the SOR parameter, the convergence of value iteration can be significantly improved.
@@
Recently, this SOR technique has been incorporated into Q-learning, resulting in a new algorithm called SOR Q-learning (SORQL) [9]. In essence, SOR Q-learning is a variant of the standard Q-learning algorithm designed to accelerate the convergence of Q-learning. The motivation for the SORQL algorithm stems from the fact that the underlying deterministic mapping of SORQL can be shown to have a contraction factor less than that of the standard Q-Bellman operator [9], thereby speeding up the convergence of the Q-learning algorithm. However, the SOR parameter in SORQL depends on the transition probability of a self-loop in the MDP, making the algorithm not entirely model-free. Furthermore, we have identified both theoretically and empirically that SOR Q-learning suffers from overestimation bias. To address these issues, in this manuscript, we propose model-free double SOR Q-learning (MF-DSORQL), which combines the idea of double Q-learning [11] with the SOR technique. The advantages of the proposed algorithm are that it is entirely model-free and has been shown, both theoretically and empirically, to have less bias than SORQL. In addition, this work provides the convergence analysis of the proposed MF-DSORQL algorithm using stochastic approximation (SA) techniques.
@@
Motivated from the work in [15], and [16], wherein the double Q-learning and SOR Q-learning in the tabular version extended to large-scale problems using the function approximation. In this manuscript, we extend the tabular version of the proposed algorithm and present a double successive over-relaxation deep Q-network (DSORDQN). The proposed algorithm is tested on roulette [11], multi-armed bandit problem [12][17], OpenAI gym’s CartPole [18], LunarLander [18] environments, Atari games and a maximization example where the state space is in the order of $10^{9}$ [19]. The experiments corroborate the theoretical findings.
I-A Related Works and Brief Literature Review
Several variants of Q-learning algorithms are available in the literature. To mention a few, in [8], speedy Q-learning (SQL) algorithm is developed to fasten the convergence of the Q-learning algorithm. At every iteration, SQL uses two successive Q-estimates in its update rule. Recently, the concept of SOR has been applied to the SQL algorithm, leading to the development of a new algorithm called generalized speedy Q-learning [20]. The DQL proposed in [11] to address the over-estimation of Q-learning was later improved in [21] using a weighted double Q-learning approach. Specifically, [21] uses a weighted combination of estimates of DQL [21]. In [12], bias-corrected Q-learning algorithm is proposed wherein a bias correction term is constructed and subtracted from the Q-learning iteration. Also, self-corrected Q-learning is proposed in [22], where a self-correcting estimator is defined to control the bias. Recently, a generalization of the Q-learning algorithm, known as maxmin Q-learning, has been proposed to address the maximisation bias, which uses $N$ estimates with $N=1$ corresponding to the Q-learning [23]. Unlike most of the algorithms mentioned above, the proposed method solves a modified Bellman’s equation. In addition, contrast to the existing works, which address the over-estimation issue of Q-learning or slower convergence of Q-learning separately, the proposed method aims to tackle both issues simultaneously.
Lately, in [24], the SORQL algorithm has been suitably modified to handle two-player zero-sum games, and consequently, a generalized minimax Q-learning algorithm was proposed. The convergence of several RL algorithms depends on the theory developed in SA [25, 26, 27, 28, 24]. Many of the iterative schemes in the literature assume the boundedness and proceed to show the convergence. For instance, in [24], the convergence of the generalized minimax Q-learning algorithm was shown to converge under the boundedness assumption. In this manuscript, we take the same approach to show the convergence of the proposed algorithms in the tabular setting.
The key contributions are summarized as follows
  • A model-based SOR Q-learning algorithm in [9] is modified into a model-free version (MF-SORQL) to make it more practically feasible.
  • The issue of overestimation in the model-based SOR Q-learning algorithm is identified and demonstrated both theoretically and empirically. Based on these findings, a new model-free algorithm, known as MF-DSORQL, is proposed.
  • The convergence of both MF-SORQL and MF-DSORQL algorithms is established using techniques from stochastic approximation theory.
  • The proposed MF-DSORQL algorithm is extended to deep RL setting and evaluated, along with its tabular version, on benchmark examples.
The paper is organized as follows: Section II covers preliminaries and convergence prerequisites. Section III presents the proposed algorithms. Section IV provides convergence analysis. Section V demonstrates numerical experiments, and Section VI concludes the paper.
@@ II Preliminaries
$r(i,a)=\sum_{j=1}^{|S|}p^{j}_{ia}r^{j}_{ia}\quad\text{and}\quad\mathcal{M}Q(j)=\max_{b\in A}Q(j,b).$ (2)
To simplify the equations, we defined the max-operator, $\max_{a\in A}Q(i,a)$ as $\mathcal{M}Q(i)$ for all $i\in S$. Also, $R_{\max}=\max_{(s,a,s^{\prime})}|r^{s^{\prime}}_{sa}|,\forall(s,a,s^{\prime})\in S\times A\times S$. These notations will be used throughout the manuscript as necessary. Note that $U$ is a contraction operator with $\gamma$ as the contraction factor, and $Q^{*}$ serves as the optimal action-value function of the operator $U$. Moreover, one can obtain an optimal policy $\pi^{*}$ from the $Q^{*}$, using the relation $\pi^{*}(i)\in\text{arg}\,\mathcal{M}Q^{*}(i),\;\forall i\in S.$
Q-learning algorithm is a SA version of Bellman’s operator $U$. Given a sample $\{i,a,j,r^{j}_{ia}\}$, current estimate $Q_{n}(i,a)$, and a suitable step-size $\beta_{n}(i,a)$, the update rule of Q-learning is as follows
$Q_{n+1}(i,a)=Q_{n}(i,a)+\beta_{n}(i,a)(r^{j}_{ia}+\gamma\,\mathcal{M}Q_{n}(j)-Q_{n}(i,a)).$ (3)
@@
$=w\bigg{(}r(i,a)+\gamma\sum_{j=1}^{|S|}p^{j}_{ia}\mathcal{M}Q(j)\bigg{)}+(1-w)\mathcal{M}Q(i)$ (4)
where $0<w\leq w^{*}$, and $w^{*}=\min\limits_{i,a}\left\{\frac{1}{1-\gamma p^{i}_{ia}}\right\}.$ The contraction factor of the map $U_{w}$ is $1-w+w\gamma$. It is interesting to note that for a suitable SOR parameter $w$, one can show that $1-w+w\gamma<\gamma$ (Lemma 4 in [9]). Consequently, the fixed-point iterative scheme corresponding to $U_{w}$ converges faster compared to the operator $U$. The SORQL algorithm is built using the operator $U_{w}$ and can be viewed as the stochastic approximation version of the fixed-point iterative scheme $Q_{n+1}=U_{w}(Q_{n})$. Therefore, one can expect that the SORQL algorithm will converge faster than that of the Q-learning. Although the fixed points of $U$ and $U_{w}$ may not be the same, the following relation between their optimal Q-values ensures that they yield the same optimal policies
$\mathcal{M}Q^{*}_{w}(i)=\mathcal{M}Q^{*}(i),\quad\forall i\in S$ (5)
@@
$\hskip 136.5733pt+(1-w)\mathcal{M}Q_{n}(i)\Big{]}$ (6)
where $0<w\leq w^{*}$, and $0\leq\beta_{n}(i,a)\leq 1$ converges to the fixed point of $U_{w}$ w.p.1 [9].
Note that the successive relaxation factor in the above equation depends on the transition probability, which makes it not entirely model-free. Owing to this and the natural concerns regarding the over-estimation of SORQL. This manuscript proposes a model-free variant of the SORQL algorithm and a model-free double SOR Q-learning algorithm.
To prove the convergence of the proposed algorithms in the tabular setting, the following result from [28] will be useful, and we conclude this section by stating it here as lemma.
@@ III-A Tabular Version
$\widetilde{P}^{j}_{ia}=\dfrac{Y_{n}[i][j][a]}{n},\quad\quad\quad n\geq 1.$ (7)
@@
$w_{n+1}=w_{n}+\alpha(n)(f(w_{n})+\epsilon_{n})\quad\quad n\geq 1$ (9)
@@
$f(w_{n})=\dfrac{1}{1-\gamma\min\limits_{i,a}p^{j}_{ia}}-w_{n},\;\epsilon_{n}=\dfrac{1}{1-\gamma\min\limits_{i,a}\widetilde{P}^{j}_{ia}}-w^{*}.$ (10)
@@
As a result of the above observations and [27], one can obtain the following theorem.
@@
7:         $b^{*}=\arg\max_{b}Q^{A}_{n}(i^{\prime}_{n},b)$, $c^{*}=\arg\max_{c}Q^{A}_{n}(i_{n},c)$
8:          $Q^{A}_{n+1}(i_{n},a_{n})=(1-\beta_{n}(i_{n},a_{n}))Q^{A}_{n}(i_{n},a_{n})+\beta_{n}(i_{n},a_{n})\big{(}w_{n}(r^{\prime}_{n}+\gamma Q^{B}_{n}(i^{\prime}_{n},b^{*}))+(1-w_{n})Q^{B}_{n}(i_{n},c^{*})\big{)}$
9:     else
10:          $d^{*}=\arg\max_{d}Q^{A}_{n}(i^{\prime}_{n},d)$, $e^{*}=\arg\max_{e}Q^{A}_{n}(i_{n},e)$
11:          $Q^{B}_{n+1}(i_{n},a_{n})=(1-\beta_{n}(i_{n},a_{n}))Q^{B}_{n}(i_{n},a_{n})+\beta_{n}(i_{n},a_{n})\big{(}w_{n}(r^{\prime}_{n}+\gamma Q^{A}_{n}(i^{\prime}_{n},d^{*}))+(1-w_{n})Q^{A}_{n}(i_{n},e^{*})\big{)}$
12:     end if
13:     $i_{n+1}=i^{\prime}_{n}$
14:end for
15:return $Q_{N-1}$
@@
If the sequence $\{w_{n}\}_{n\geq 0}$ in Algorithm 1 is a constant sequence, with $w$ as the constant and $0<w\leq w^{*}$, we refer to this algorithm as double SOR Q-learning (DSORQL). Therefore, given the samples $\{i,a,j,r^{j}_{ia}\}$ the update rule for the DSORQL is as follows
With probability $0.5$ update $\widetilde{Q}^{A}$
$\widetilde{Q}^{A}_{n+1}(i,a)$ $=(1-\beta_{n})\widetilde{Q}^{A}_{n}(i,a)$ (12)
@@
else update $\widetilde{Q}^{B}$
$\widetilde{Q}^{B}_{n+1}(i,a)$ $=(1-\beta_{n})\widetilde{Q}^{B}_{n}(i,a)$ (13)
$\;+\beta_{n}\Big{[}w\big{(}r^{j}_{ia}+\gamma\widetilde{Q}^{A}_{n}(j,d^{*})\big{)}+(1-w)\big{(}\widetilde{Q}^{A}_{n}(i,e^{*})\big{)}\Big{]}$
@@ III-B Deep RL Version
The proposed double SOR deep Q-network (DSORDQN) is a deep RL version of the tabular DSORQL. DSORDQN utilizes the technique of SOR method with DDQN. Specifically, similar to DDQN we use online Q-network $Q(S,a;\theta)$ for selecting the actions, and the target Q-network $Q(S,a;\theta^{\prime})$, to evaluate the actions. The gradient descent will be performed on the following loss function: $L(\theta_{i})=\mathbb{E}\big{[}(y_{j}-Q(s,a;\theta_{i}))^{2}\big{]},$ where $y_{j}=w(r(i_{j},a_{j})+\gamma Q(i_{j+1},\operatorname*{arg\,max}_{b}Q(i_{j+1},b;\theta_{i});\theta^{\prime}_{i})+(1-w)Q(i_{j},\operatorname*{arg\,max}_{e}Q(i_{j},e;\theta_{i});\theta^{\prime}_{i})$ for sample $(i_{j},a_{j},r(i_{j},a_{j}),i_{j+1})$ from the replay buffer with $w>1$.
Interestingly, the advantages of DDQN and SORDQN transfer to the proposed DSORDQN algorithm, resulting in more stable and reliable learning at scale.